# Universal contributions to charge fluctuations in spin chains at finite temperature

### 17 Jan 2024 v1

@@ I Introduction
At finite temperature, the state of a quantum many-body system with $\mathrm{U}(1)$ symmetry is conveniently described by the thermal density matrix $\rho=\frac{1}{Z}\,\mathrm{e}^{-\beta(H-\mu Q)}$ in the grand canonical ensemble, where $\beta$ is the inverse temperature and $\mu$ is the chemical potential. It is well-known that even though the total charge $Q$ is not conserved in the ensemble, it has a well-defined average with fluctuations suppressed in the thermodynamic limit. Higher moments of $Q$ are also suppressed as the distribution approaches a Gaussian one. Our main objective in this work is to show that the charge fluctuations may contain universal information about the system, especially when the $\mathrm{U}(1)$ charge-conservation symmetry exhibits certain types of ’t Hooft anomaly (usually a mixed anomaly with other global symmetries). To extract this information, it is most useful to consider the following generating function (also known as the full counting statistics Belzig and Nazarov (2001); Levitov and Reznikov (2004); Song et al. (2012, 2010))
@@
Another motivation to study $G(\theta)$ comes from its connection with the disorder parameter in the (2+1)d ground state of a gapped Hamiltonian Wang et al. (2021); Wu et al. (2021); Chen et al. (2022); Cuomo et al. (2022). Suppose $A$ is a subregion of the (2+1)d system, and denote by $Q_{A}$ the total charge inside $A$. We define the disorder parameter as $\langle\,\mathrm{e}^{\mathrm{i}\theta Q_{A}}\rangle$, where the expectation value now is taken with respect to the ground state. When the ground state is gapped and preserves the $\mathrm{U}(1)$ symmetry, one expects for a region $A$ of size $L_{A}$
@@
While the ’t Hooft anomalies in these two systems appear quite different, they share a common feature: a $\pi$ flux of the U(1) symmetry carries a projective representation protected by the other symmetries. This type of ’t Hooft anomaly is often referred to as the “type-III” anomaly in literature de Wild Propitius (1995). The type-III structure of the anomaly turns out to be crucial for the quantization of $\gamma(\pi)$.
@@
More generally, we provide arguments for the quantization of $\gamma(\pi)$ based on the matrix product representation of the density matrix for both cases mentioned above, generalizing an argument in Zang et al. . As a by-product, we find that in both cases the U(1) symmetry can be reduced to the $\mathds{Z}_{2}$ subgroup without affecting the results, since the type-III anomaly structure is preserved.
@@ II Charge fluctuations in a CFT
Here $d_{\theta}$ is the ground state degeneracy in the defect sector. $h_{\theta}$ is the scaling dimension of the $\mathrm{U}(1)$ defect. For a rational CFT, the presence of a $\mathrm{U}(1)$ symmetry implies that the CFT has $\mathrm{U}(1)_{k}\times\mathrm{U}(1)_{-{k}}$ Kac-Moody algebra, with $k$ being the level, and for $\theta\in[0,\pi]$, $h_{\theta}$ is given by the charged Cardy formula Dyer et al. (2018); Hosseini et al. (2020); Fan et al. (2023):
@@
We note that the argument can be easily generalized to other symmetries. In fact, Eq. (12) holds without any modification, except that $h_{\theta}$ should be replaced by the conformal weight of the corresponding defect operator. Essentially equivalent result was found in the context of topological disorder parameter in a (2+1)d gapped state in Chen et al. (2022).
@@
As expected on general grounds, $\alpha(\theta)$ is “local” and insensitive to global boundary condition. Notice that $\alpha(\theta)$ exhibits a cusp at $\theta=\pi$, which appears to be a general feature of CFTs with a U(1) symmetry. A similar cusp was found in the coefficient of the leading term in the $\mathrm{U}(1)$ disorder parameter in a (1+1)d Luttinger liquid or a (2+1)d Fermi liquid Tan and Ryu (2020); Jiang et al. (2023); Cai and Cheng .
@@
However, an important caveat in this argument for the robustness of $\gamma$ and $\omega$ is the dependence of $\eta_{w}$ on the microscopic physics. It has been understood now that certain microscopic conditions, such as the filling factor, enter the low-energy theory as background topological defects required by anomaly matching Cheng et al. (2016); Metlitski and Thorngren (2018); Cho et al. (2017); Jian et al. (2018). As an example, consider a $c=1$ Luttinger liquid in a lattice system with $\mathrm{U}(1)$ filling factor $\nu$ (i.e. the average charge per unit cell is $\nu$). The low-energy physics of a system with periodic boundary condition of $N$ unit cells should be described by the Luttinger liquid with a defect $\eta_{w}=-N\nu$ Cheng and Seiberg (2023); Metlitski and Thorngren (2018); Else et al. (2021). Since the filling can be continuously tuned by applying a chemical potential, the defect parameter and thus $\gamma(\pi)$ can also be changed continuously.
@@ III Spin-1/2 chain with LSM anomaly
Notice that the O(2) symmetry is on-site, so it is not anomalous on its own. However, because each spin-1/2 site carries a projective representation of the O(2) symmetry, the system exhibits a LSM anomaly between the O(2) and the lattice translation Cheng et al. (2016); Jian et al. (2018); Metlitski and Thorngren (2018); Else and Thorngren (2020).
Without the $\mathds{Z}_{2}$ symmetry, there is no ’t Hooft anomaly between the U(1) and lattice translation symmetry. However, if the total filling is fixed, e.g. $Q=0$ (corresponding to half filling in the hard-core boson basis), the system exhibits the “filling anomaly” for $\mathrm{U}(1)$ and lattice translation Oshikawa (2000); Cheng et al. (2016); Cho et al. (2017); Jian et al. (2018); Cheng and Seiberg (2023), forbidding the existence of a symmetric trivial state. However, unlike the O(2) LSM anomaly, the filling anomaly only appears in the subspace constrained to have a fixed filling. Therefore, in a grand canonical ensemble, where the total charge is allowed to fluctuate, the filling anomaly is absent.
@@
$Z_{00}(\theta)\approx\,\mathrm{e}^{\frac{N}{2\pi}I_{1}\left(\theta,\beta\right)},$ (27)
where the integral $I_{1}(\theta,\beta)$ is defined in (104). Here $\approx$ means that there are $\mathcal{O}(N^{-1})$ corrections in the exponential.
@@
$Z_{11}(0)=\prod_{j=\frac{1}{2}}^{N-\frac{1}{2}}(1-\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}})$ $\approx 4\,\mathrm{e}^{\frac{N}{2\pi}I_{2}(\beta)}.$ (28)
@@
It is worth noting that these results are valid for any temperature $\beta\neq 0$. For $\beta=0$, we instead get $G(\theta)=\left(\cos\frac{\theta}{2}\right)^{N}$, which still gives $\gamma(\theta)=0$ for any $\theta\neq\pi$, but $G(\pi)=0$ so $\gamma(\pi)$ is not well-defined.
@@
so $\gamma(\theta)=0$. The phase factor of $G(\theta)$ now depends on $\beta$ and $\nu$ in a complicated way. Only when $\beta$ is large (i.e. low temperature), the phase factor approaches the CFT form $\,\mathrm{e}^{\mathrm{i}N\nu[\theta]_{\pi}}$.
For $\theta=\pi$ we are able to determine the phase factor as well:
@@ III.1 Quantization of $\gamma(\pi)$
To proceed we follow the argument presented in Zang et al. . First we represent a translation-invariant density matrix as a matrix product operator (MPO) Verstraete et al. (2004); Perez-Garcia et al. (2007); Cirac et al. (2017, 2021):
@@
It is known that such an approximation is always possible for thermal states of local Hamiltonians Hastings (2006); Molnar et al. (2015); Kuwahara et al. (2021); Huang (2021). Furthermore, since there is no long-range order at finite temperature in 1D, we expect that the MPO is injective Perez-Garcia et al. (2007); Cirac et al. (2021), i.e. the corresponding transfer matrix has a non-degenerate leading eigenvalue.
@@
In other words, $W_{X}$ and ${M}_{Z}$ anticommute. Therefore, all eigenvalues of ${M}_{Z}$ must come in pairs ±$\pm\lambda$, including those with the largest magnitude. In principle, it is possible that $\left\lvert\lambda\right\rvert$ has degeneracy more than 2. However, given that the leading eigenvalue of $M$ is non-degenerate, we expect that $M_{Z}$ has no more degeneracy in the leading $|\lambda|$ than what is required by the symmetry condition in Eq. (40). That is, the leading eigenvalues of ${M}_{Z}$ should be ±$\pm\lambda_{\rm max}$  [^1] It is possible that $\lambda_{\rm max}=0$, in which case $\gamma$ becomes meaningless. This is what happens at $\beta=0$. With this assumption,
@@
Notice that the result applies to any system with $\mathds{Z}_{2}\times\mathds{Z}_{2}$ LSM anomaly Cheng et al. (2016); Po et al. (2017); Ogata and Tasaki (2019); Ogata et al. (2021), since the only symmetry property used here is $X_{n}Z_{n}=-Z_{n}X_{n}$. In fact, the only assumptions needed on $\rho$ are translation invariance and the $X$ symmetry.
@@
can be made hermitian for $\Delta\geq 0$ Tang et al. (2020). Thus Eq. (46) applies to $H_{\rm XXZ}$ as well.
@@ IV Spin chain with mixed $\mathrm{O}(2)\times\mathds{Z}_{2}$ anomaly
Now we discuss the ’t Hooft anomaly of the internal $\mathds{Z}_{2}^{X}\times\mathrm{O}(2)$ symmetry. As alluded to above, there is a mixed anomaly between $\mathds{Z}_{2}^{X}$ and O(2). In fact, the same kind of anomaly is already present when restricted to $\mathds{Z}_{2}\subset\mathrm{U}(1)$. The $\mathds{Z}_{2}\times\mathds{Z}_{2}^{C}\times\mathds{Z}_{2}^{X}$ symmetry has an anomaly associated with the so-called “type-III” 3-cocycle de Wild Propitius (1995). Physically, it is characterized by the $\mathds{Z}_{2}$ defect transforming projectively under $\mathds{Z}_{2}^{\cal C}\times\mathds{Z}_{2}^{X}$.
Moreover, even the $\mathds{Z}_{2}^{C}$ symmetry is ignored, the remaining $\mathrm{U}(1)\times\mathds{Z}_{2}^{X}$ symmetry is still anomalous. The system can be viewed as a lattice model for the edge of a bosonic SPT state protected by the $\mathrm{U}(1)\times\mathds{Z}_{2}$ symmetry Chen et al. (2013); Metlitski (2019); Son and Alicea (2019); Horinouchi (2020). Turning on a non-zero chemical potential breaks $\mathds{Z}_{2}^{C}$ but preserves $\mathds{Z}_{2}^{X}$.
@@
which was first considered in Levin and Gu (2012) as an edge model of the nontrivial (2+1)d $\mathds{Z}_{2}$ SPT state. The model is exactly solvable, and can be mapped to a gauged XX spin chain as follows. Introduce a dual representation of the system Levin and Gu (2012):
@@
To understand the physical meaning of $\gamma$ and $\omega$, let us consider two limits. First, at low temperature $\beta\rightarrow\infty$, the system is described by a $c=1$ Luttinger liquid with $R=\frac{1}{\sqrt{2}}$ Cheng and Seiberg (2023). The $\mathrm{U}(1)$ charge $Q$ is identified with $Q_{m}$ in the low energy theory, and $X$ becomes $\,\mathrm{e}^{\mathrm{i}\pi Q_{w}}$. In addition, the lattice unit translation leads to an emanant $\mathds{Z}_{4}$ symmetry $\,\mathrm{e}^{-\frac{\mathrm{i}\pi}{2}Q_{w}}$ Cheng and Seiberg (2023). Namely, the lattice translation $T$ has the following representation in the low-energy theory:
@@ V Conclusions and Discussions
Following this line of thought, one expects that similar results should hold for fermionic systems. For example, in a (1+1)d fermionic system with $\mathds{Z}_{2}\times\mathds{Z}_{2}^{F}$ symmetry ($\mathds{Z}_{2}^{F}$ stands for the fermion parity conservation), the ’t Hooft anomaly is classified by $\mathds{Z}_{8}$. The generator of the $\mathds{Z}_{8}$ is characterized by the $\mathds{Z}_{2}$ symmetry defect carrying a Majorana zero mode. Thus we expect that if one measures the expectation value of the total $\mathds{Z}_{2}$ charge in a thermal state, $\gamma$ should take a universal value $\gamma=\ln\sqrt{2}$. Another system with a mathematically similar anomaly is a translation-invariant chain of Majorana modes Cheng (2019), where the lattice translation has a mixed anomaly with fermion parity. In this case, we expect that the thermal expectation value of the translation operator contains a universal correction $\ln\sqrt{2}$.

### 03 Mar 2024 v2

@@ I Introduction
At finite temperature, the state of a quantum many-body system with $\mathrm{U}(1)$ symmetry is conveniently described by the thermal density matrix $\rho=\frac{1}{Z}\,\mathrm{e}^{-\beta(H-\mu Q)}$ in the grand canonical ensemble, where $\beta$ is the inverse temperature and $\mu$ is the chemical potential. It is well-known that even though the total charge $Q$ is not conserved in the ensemble, it has a well-defined average with fluctuations suppressed in the thermodynamic limit. Higher moments of $Q$ are also suppressed as the distribution approaches a Gaussian one. Our main objective in this work is to show that the charge fluctuations may contain universal information about the system, especially when the $\mathrm{U}(1)$ charge-conservation symmetry exhibits certain types of ’t Hooft anomaly (usually a mixed anomaly with other global symmetries). To extract this information, it is most useful to consider the following generating function (also known as the full counting statistics [1, 2, 3, 4])
@@
Another motivation to study $G(\theta)$ comes from its connection with the disorder parameter in the (2+1)d ground state of a gapped Hamiltonian [5, 6, 7, 8]. Suppose $A$ is a subregion of the (2+1)d system, and denote by $Q_{A}$ the total charge inside $A$. We define the disorder parameter as $\langle\,\mathrm{e}^{\mathrm{i}\theta Q_{A}}\rangle$, where the expectation value now is taken with respect to the ground state. When the ground state is gapped and preserves the $\mathrm{U}(1)$ symmetry, one expects for a region $A$ of size $L_{A}$
@@
While the ’t Hooft anomalies in these two systems appear quite different, they share a common feature: a $\pi$ flux of the U(1) symmetry carries a projective representation protected by the other symmetries. This type of ’t Hooft anomaly is often referred to as the “type-III” anomaly in literature [9]. The type-III structure of the anomaly turns out to be crucial for the quantization of $\gamma(\pi)$.
@@
More generally, we provide arguments for the quantization of $\gamma(\pi)$ based on the matrix product representation of the density matrix for both cases mentioned above, generalizing an argument in [10]. As a by-product, we find that in both cases the U(1) symmetry can be reduced to the $\mathds{Z}_{2}$ subgroup without affecting the results, since the type-III anomaly structure is preserved.
@@ II Charge fluctuations in a CFT
Here $d_{\theta}$ is the ground state degeneracy in the defect sector. $h_{\theta}$ is the scaling dimension of the $\mathrm{U}(1)$ defect. For a rational CFT, the presence of a $\mathrm{U}(1)$ symmetry implies that the CFT has $\mathrm{U}(1)_{k}\times\mathrm{U}(1)_{-{k}}$ Kac-Moody algebra, with $k$ being the level, and for $\theta\in[0,\pi]$, $h_{\theta}$ is given by the charged Cardy formula [11, 12, 13]:
@@
We note that the argument can be easily generalized to other symmetries. In fact, Eq. (12) holds without any modification, except that $h_{\theta}$ should be replaced by the conformal weight of the corresponding defect operator. Essentially equivalent result was found in the context of topological disorder parameter in a (2+1)d gapped state in [7].
@@
As expected on general grounds, $\alpha(\theta)$ is “local” and insensitive to global boundary condition. Notice that $\alpha(\theta)$ exhibits a cusp at $\theta=\pi$, which appears to be a general feature of CFTs with a U(1) symmetry. A similar cusp was found in the coefficient of the leading term in the $\mathrm{U}(1)$ disorder parameter in a (1+1)d Luttinger liquid or a (2+1)d Fermi liquid [14, 15, 16].
@@
However, an important caveat in this argument for the robustness of $\gamma(\theta)$ and $\omega(\theta)$ is the dependence of $\eta_{w}$ on the microscopic physics. It has been understood now that certain microscopic conditions, such as the filling factor, enter the low-energy theory as background topological defects required by anomaly matching [17, 18, 19, 20]. As an example, consider a $c=1$ Luttinger liquid in a lattice system with $\mathrm{U}(1)$ filling factor $\nu$ (i.e. the average charge per unit cell is $\nu$). The low-energy physics of a system with periodic boundary condition of $N$ unit cells should be described by the Luttinger liquid with a defect $\eta_{w}=-N\nu$ [21, 18, 22]. Since the filling can be continuously tuned by applying a chemical potential, the defect parameter and thus $\gamma(\pi)$ can also be changed continuously.
@@ III Spin-1/2 chain with LSM anomaly
Notice that the O(2) symmetry is on-site, so it is not anomalous on its own. However, because each spin-1/2 site carries a projective representation of the O(2) symmetry, the system exhibits a LSM anomaly between the O(2) and the lattice translation [17, 20, 18, 23].
Without the $\mathds{Z}_{2}$ symmetry, there is no ’t Hooft anomaly between the U(1) and lattice translation symmetry. However, if the total filling is fixed, e.g. $Q=0$ (corresponding to half filling in the hard-core boson basis), the system exhibits the “filling anomaly” for $\mathrm{U}(1)$ and lattice translation [24, 17, 19, 20, 21], forbidding the existence of a symmetric trivial state. However, unlike the O(2) LSM anomaly, the filling anomaly only appears in the subspace constrained to have a fixed filling. Therefore, in a grand canonical ensemble, where the total charge is allowed to fluctuate, the filling anomaly is absent.
@@
$Z_{00}(\theta)\approx\,\mathrm{e}^{\frac{N}{2\pi}I\left(\theta,\beta\right)},$ (27)
where the integral $I(\theta,\beta)$ is defined in (105). Here $\approx$ means that there are $\mathcal{O}(N^{-1})$ corrections in the exponential.
@@
$Z_{11}(0)=\prod_{j=\frac{1}{2}}^{N-\frac{1}{2}}(1-\,\mathrm{e}^{4\beta\cos\frac{2\pi j}{N}})$ $\approx 4\,\mathrm{e}^{\frac{N}{2\pi}\Re I(\pi,\beta)}.$ (28)
@@
It is worth noting that these results are valid for any temperature $\beta\neq 0$. For $\beta=0$, we instead get $G(\theta)=\left(\cos\frac{\theta}{2}\right)^{N}$, which still gives $\gamma(\theta)=0$ for any $\theta\neq\pi$, but $G(\pi)=0$, so $\gamma(\pi)$ is not well-defined.
@@
so $\gamma(\theta)=0$. We plot $\alpha(\theta)$ for several different temperatures in Fig. 2. They agree very well with the CFT result, and exhibit a cusp at $\theta=\pi$ for all temperatures. The cusp becomes sharper as the temperature increases.
The phase factor $\omega(\theta)$ now depends on $\beta$ and $\nu$ in a complicated way. This is expected, since for small $\theta$, one has
$\omega(\theta)\approx\,\mathrm{e}^{\mathrm{i}N\nu(\beta)\theta},$ (34)
where $\nu(\beta)$ is the average filling at temperature $\beta$. Only when $\beta$ is large (i.e. low temperature), the phase factor approaches the CFT form $\,\mathrm{e}^{\mathrm{i}N\nu[\theta]_{\pi}}$. We plot $\arg(\omega)$ for the $\nu=\frac{1}{10}$ case in Fig 1.
Figure 1: $\frac{1}{N}\arg\omega(\theta)$ for $\nu=\frac{1}{10}$. At low temperature, the numerical results agree with the CFT value, while at high temperture, they approach $\frac{1}{2}[\theta]_{\pi}$. For $\nu=\frac{1}{2}$, all the lines collapse.
Figure 2: $\alpha(\theta)$ for $\nu=\frac{2}{3}$ at different $\beta$.
For $\theta=\pi$, we are able to determine the phase factor as well:
@@ III.1 Quantization of $\gamma(\pi)$
To proceed we follow the argument presented in [10]. First we represent a translation-invariant density matrix as a matrix product operator (MPO) [25, 26, 27, 28]:
@@
It is known that such an approximation is always possible for thermal states of local Hamiltonians [29, 30, 31, 32]. Furthermore, since there is no long-range order at finite temperature in 1D, we expect that the MPO is injective [26, 28], i.e. the corresponding transfer matrix has a non-degenerate leading eigenvalue.
@@
In other words, $W_{X}$ and ${M}_{Z}$ anticommute. Therefore, all eigenvalues of ${M}_{Z}$ must come in pairs ±$\pm\lambda$, including those with the largest magnitude. In principle, it is possible that $\left\lvert\lambda\right\rvert$ has degeneracy more than 2. However, given that the leading eigenvalue of $M$ is non-degenerate, we expect that $M_{Z}$ has no more degeneracy in the leading $|\lambda|$ than what is required by the symmetry condition in Eq. (41). That is, the leading eigenvalues of ${M}_{Z}$ should be ±$\pm\lambda_{\rm max}$   [^1] It is possible that $\lambda_{\rm max}=0$, in which case $\gamma$ becomes meaningless. This is what happens at $\beta=0$.. With this assumption,
@@
Notice that the result applies to any system with $\mathds{Z}_{2}\times\mathds{Z}_{2}$ LSM anomaly [17, 34, 35, 36], since the only symmetry property used here is $X_{n}Z_{n}=-Z_{n}X_{n}$. In fact, the only assumptions needed on $\rho$ are translation invariance and the $X$ symmetry.
@@
can be made hermitian for $\Delta\geq 0$ [37]. Thus Eq. (47) applies to $H_{\rm XXZ}$ as well.
@@ IV Spin chain with mixed $\mathrm{O}(2)\times\mathds{Z}_{2}$ anomaly
Now we discuss the ’t Hooft anomaly of the internal $\mathds{Z}_{2}^{X}\times\mathrm{O}(2)$ symmetry. As alluded to above, there is a mixed anomaly between $\mathds{Z}_{2}^{X}$ and O(2). In fact, the same kind of anomaly is already present when restricted to $\mathds{Z}_{2}\subset\mathrm{U}(1)$. The $\mathds{Z}_{2}\times\mathds{Z}_{2}^{C}\times\mathds{Z}_{2}^{X}$ symmetry has an anomaly associated with the so-called “type-III” 3-cocycle [9]. Physically, it is characterized by the $\mathds{Z}_{2}$ defect transforming projectively under $\mathds{Z}_{2}^{\cal C}\times\mathds{Z}_{2}^{X}$.
Moreover, even the $\mathds{Z}_{2}^{C}$ symmetry is ignored, the remaining $\mathrm{U}(1)\times\mathds{Z}_{2}^{X}$ symmetry is still anomalous. The system can be viewed as a lattice model for the edge of a bosonic SPT state protected by the $\mathrm{U}(1)\times\mathds{Z}_{2}$ symmetry [38, 39, 40, 41]. Turning on a non-zero chemical potential breaks $\mathds{Z}_{2}^{C}$ but preserves $\mathds{Z}_{2}^{X}$.
@@
which was first considered in [42] as an edge model of the nontrivial (2+1)d $\mathds{Z}_{2}$ SPT state. The model is exactly solvable, and can be mapped to a gauged XX spin chain as follows. Introduce a dual representation of the system [42]:
@@
To understand the physical meaning of $\gamma$ and $\omega$, let us consider two limits. First, at low temperature $\beta\rightarrow\infty$, the system is described by a $c=1$ Luttinger liquid with $R=\frac{1}{\sqrt{2}}$ [21]. The $\mathrm{U}(1)$ charge $Q$ is identified with $Q_{m}$ in the low energy theory, and $X$ becomes $\,\mathrm{e}^{\mathrm{i}\pi Q_{w}}$. In addition, the lattice unit translation leads to an emanant $\mathds{Z}_{4}$ symmetry $\,\mathrm{e}^{-\frac{\mathrm{i}\pi}{2}Q_{w}}$ [21]. Namely, the lattice translation $T$ has the following representation in the low-energy theory:
@@ V Conclusions and Discussions
Following this line of thought, one expects that similar results should hold for fermionic systems. For example, in a (1+1)d fermionic system with $\mathds{Z}_{2}\times\mathds{Z}_{2}^{F}$ symmetry ($\mathds{Z}_{2}^{F}$ stands for the fermion parity conservation), the ’t Hooft anomaly is classified by $\mathds{Z}_{8}$. The generator of the $\mathds{Z}_{8}$ is characterized by the $\mathds{Z}_{2}$ symmetry defect carrying a Majorana zero mode. Thus we expect that if one measures the expectation value of the total $\mathds{Z}_{2}$ charge in a thermal state, $\gamma$ should take a universal value $\gamma=\ln\sqrt{2}$. Another system with a mathematically similar anomaly is a translation-invariant chain of Majorana modes [43], where the lattice translation has a mixed anomaly with fermion parity. In this case, we expect that the thermal expectation value of the translation operator contains a universal correction $\ln\sqrt{2}$.